OpenStudy (anonymous):

Are y=-1/2x+3/2 and 5x-10y=15 parallel? Please explain.

OpenStudy (shamil98):

parallel lines have equal slopes. 5x - 10y = 15 the slope of this is m = - a/b m = -5/-10 m = 1/2 the slope of y = -1/2x + 3/2 is - 1/2 so no the equations do not have equal slopes, and therefore are not parallel